D. Bag of mice
time limit per test

2 seconds

memory limit per test

256 megabytes

input

standard input

output

standard output

The dragon and the princess are arguing about what to do on the New Year's Eve. The dragon suggests flying to the mountains to watch fairies dancing in the moonlight, while the princess thinks they should just go to bed early. They are desperate to come to an amicable agreement, so they decide to leave this up to chance.

They take turns drawing a mouse from a bag which initially contains w white and b black mice. The person who is the first to draw a white mouse wins. After each mouse drawn by the dragon the rest of mice in the bag panic, and one of them jumps out of the bag itself (the princess draws her mice carefully and doesn't scare other mice). Princess draws first. What is the probability of the princess winning?

If there are no more mice in the bag and nobody has drawn a white mouse, the dragon wins. Mice which jump out of the bag themselves are not considered to be drawn (do not define the winner). Once a mouse has left the bag, it never returns to it. Every mouse is drawn from the bag with the same probability as every other one, and every mouse jumps out of the bag with the same probability as every other one.

Input

The only line of input data contains two integers w and b (0 ≤ w, b ≤ 1000).

Output

Output the probability of the princess winning. The answer is considered to be correct if its absolute or relative error does not exceed10 - 9.

Examples
input
1 3
output
0.500000000
input
5 5
output
0.658730159
Note

Let's go through the first sample. The probability of the princess drawing a white mouse on her first turn and winning right away is 1/4. The probability of the dragon drawing a black mouse and not winning on his first turn is 3/4 * 2/3 = 1/2. After this there are two mice left in the bag — one black and one white; one of them jumps out, and the other is drawn by the princess on her second turn. If the princess' mouse is white, she wins (probability is 1/2 * 1/2 = 1/4), otherwise nobody gets the white mouse, so according to the rule the dragon wins.

大佬题解:http://blog.csdn.net/swust_Three/article/details/68941926

#include<cstdio>
using namespace std;
double dp[][];
int main()
{
int w,b;
scanf("%d%d",&w,&b);
for(int i=;i<=w;i++) dp[i][]=1.0;
for(int i=;i<=w;i++)
for(int j=;j<=b;j++)
{
dp[i][j]=1.0*i/(i+j);
if(j>=) dp[i][j]+=1.0*j/(i+j)*(j-)/(i+j-)*(j-)/(i+j-)*dp[i][j-];
if(j>=) dp[i][j]+=1.0*j/(i+j)*(j-)/(i+j-)*i/(i+j-)*dp[i-][j-];
}
printf("%.9lf",dp[w][b]);
}

Codeforces 148 D Bag of mice的更多相关文章

  1. CodeForces:148D-D.Bag of mice

    Bag of mice time limit per test 2 seconds memory limit per test 256 megabytes Program Description Th ...

  2. 【codeforces 148D】 Bag of mice

    http://codeforces.com/problemset/problem/148/D (题目链接) 题意 包中有w个白鼠,b个黑鼠.公主和龙轮流画老鼠,公主先画,谁先画到白鼠谁就赢.龙每画完一 ...

  3. 【Codeforces 105D】 Bag of mice

    [题目链接] http://codeforces.com/contest/148/problem/D [算法] 概率DP f[w][b]表示还剩w只白老鼠,b只黑老鼠,公主胜利的概率,那么 : 1. ...

  4. Codeforces Round #105 (Div. 2) D. Bag of mice 概率dp

    题目链接: http://codeforces.com/problemset/problem/148/D D. Bag of mice time limit per test2 secondsmemo ...

  5. Codeforces 148D 一袋老鼠 Bag of mice | 概率DP 水题

    除非特别忙,我接下来会尽可能翻译我做的每道CF题的题面! Codeforces 148D 一袋老鼠 Bag of mice | 概率DP 水题 题面 胡小兔和司公子都认为对方是垃圾. 为了决出谁才是垃 ...

  6. Bag of mice(CodeForces 148D )

    D. Bag of mice time limit per test 2 seconds memory limit per test 256 megabytes input standard inpu ...

  7. Bag of mice(概率DP)

    Bag of mice  CodeForces - 148D The dragon and the princess are arguing about what to do on the New Y ...

  8. CF 148D. Bag of mice (可能性DP)

    D. Bag of mice time limit per test 2 seconds memory limit per test 256 megabytes input standard inpu ...

  9. 【CF148D】 Bag of mice (概率DP)

    D. Bag of mice time limit per test 2 seconds memory limit per test 256 megabytes input standard inpu ...

随机推荐

  1. mac安装php分词工具xunsearch出现找不到bio.h的解决办法

    下载xunsearch后安装出现如下错误,在xunsearch官方论坛未找到答案,此方案不仅用于参考解决安装xunsearch,其它编辑安装出现的问题同样可以参考 -n Checking scws . ...

  2. 01—为什么使用java

    Java解决的问题 1.指针问题 java里面没有指针,用引用解决指针问题,但是引用是一种限制的指针,不能参与整数运行和指向任意位置的内存,并且不用显示回收对象 引用地址:http://blog.cs ...

  3. input value="值栈的值"

    <input type="text" value="<s:property value="myp.begintime"/>" ...

  4. CentOS 7 网卡命名修改为eth0格式

    Linux 操作系统的网卡设备的传统命名方式是 eth0.eth1.eth2等,而 CentOS7 提供了不同的命名规则,默认是基于固件.拓扑.位置信息来分配.这样做的优点是命名全自动的.可预知的,缺 ...

  5. #Leetcode# 922. Sort Array By Parity II

    https://leetcode.com/problems/sort-array-by-parity-ii/ Given an array A of non-negative integers, ha ...

  6. Hive查看执行日志

    HIVE-如何查看执行日志 HIVE既然是运行在hadoop上,最后又被翻译为MapReduce程序,通过yarn来执行.所以我们如果想解决HIVE中出现的错误,需要分成几个过程 HIVE自身翻译成为 ...

  7. SSH框架配置

    --------------------------------applicationContext.xml-------------------------------- <?xml vers ...

  8. 小程序 switch按钮

    <view class='pay-switch'> <switch color='#1F3238' data-gongprice='{{gongprice}}' data-disco ...

  9. 【Python】Python基础教程系列目录

    Python是一个高层次的结合了解释性.编译性.互动性和面向对象的脚本语言. 在现在的工作及开发当中,Python的使用越来越广泛,为了方便大家的学习,Linux大学 特推出了 <Python基 ...

  10. js中关于array的常用方法

    最近总结了一些关于array中的常用方法, 其中大部分的方法来自于<JavaScript框架设计>这本书, 如果有更好的方法,或者有关于string的别的常用的方法,希望大家不吝赐教. 第 ...